# Enhancing LLM Code Generation with Ensembles: A Similarity-Based Selection Approach

### v1

@@ 4.2.1 RQ1: How does EnsLLM perform compared to standalone LLMs?
To investigate the failed cases of the HumanEval benchmark, where we found that in 121 out of 164 cases, at least two models generated correct answers, and in all such cases, EnsLLM successfully selected the correct program. This supports our assumption that in our similarity-based ensemble approach, correct programs complement each other, reinforcing their selection. Among the remaining problems, in 28 cases, at least one model generated the correct program, and EnsLLM selected 27 of them correctly. The one case where EnsLLM failed was due to three LLMs generating similar mistakes in the program, further supporting our assumption that independent LLMs are less prone to making identical mistakes.
@@ 4.3 Threats to Validity
For external validity, our evaluation is limited to HumanEval and LiveCodeBench, which may not fully represent complex, real-world programming challenges. However, these two benchmarks are widely recognized in code generation research and have been used in several similar studies to evaluate the functional correctness of LLM-generated programs. Their extensive use ensures a standardized comparison with existing approaches, mitigating the threat to generalizability.
@@ 5 Related Work
LLMs have also been explored in broader software engineering tasks, including automated bug detection, test case generation, code summarization, and security vulnerability analysis [34, 29]. For instance, they can assist in writing unit tests by generating test cases from function signatures, improving test coverage while reducing developer effort [29]. Additionally, LLMs are being integrated into code refactoring tools, where they suggest optimizations and enhance code readability [44]. Beyond these applications, LLMs are transforming software engineering workflows, influencing how software is developed [19, 27], enhancing developer productivity [13, 45], and assisting in security analysis [32]. However, challenges persist in ensuring correctness, maintainability, and security, requiring further refinement and validation before LLM-generated code can be reliably integrated into real-world software systems.
The mixture of experts (MoE) framework [21] combines multiple specialized models with a gating network to select the best expert for each input, using Expectation-Maximization for training. Fedus et al. [16] scaled MoE to trillion-parameter Switch Transformers with sparse activation for efficient language processing.
Ensemble learning has been widely used to improve robustness and accuracy in software engineering tasks. Studies show that combining multiple LLM-generated solutions through voting, weighted averaging, or meta-learning can enhance correctness [43, 37]. Some approaches leverage multi-model ensembles, where different LLMs specialize in different aspects of code quality, leading to higher functional correctness [37]. Despite these advancements, computational efficiency remains a challenge when using multiple models for selection and refinement. Also, EnsLLM is the first approach that ensembles multiple LLMs for code generation.

### v2

@@ 4.2.3 RQ3: How much accuracy can be gained using only free LLMs?
5 Discussions
5.1 Qualitative Analysis
To better understand the behavior of EnsLLM  we conducted a qualitative analysis of the failed cases on the HumanEval benchmark. Out of 164 problems, we found that in 131 cases, at least two individual models independently generated correct solutions. In all of these instances, EnsLLM successfully selected the correct program. This strongly supports our hypothesis that correct outputs reinforce one another in our similarity-based ensemble, making them more likely to be chosen.
For 18 problems (out of the rest 33 problems), only a single model produced the correct solution for each of those problems. Even in these more challenging scenarios, EnsLLM correctly identified the correct program in 17 cases, as no incorrect programs made similar mistakes that would have reinforced one another. The one failure occurred when three models (Llama 3.2, Gemma 2, and CodeLlama) produced incorrect programs with similar mistakes, leading EnsLLM to select the wrong output. This illustrates a key limitation of the approach: when multiple models converge on similar errors, the ensemble may amplify rather than avoid them. However, such failures were rare and further support the assumption that diverse model behaviors reduce correlated mistakes. For the remaining 15 cases, no correct solutions were available for EnsLLM to select.
5.2 Scalability Analysis
Collecting outputs sequentially from each of the LLMs takes 94.22 seconds per problem on average, with an associated cost of approximately 0.005 USD per proprietary model. While parallel execution can significantly reduce latency, running all 14 models, especially at scale, poses practical challenges due to resource demands and infrastructure constraints. The ensemble evaluation stage adds further overhead: computing all 91 pairwise similarity scores using CodeBLEU takes 86.21 seconds per problem, and running behavioral checks with CrossHair requires 212.38 seconds. These components are critical to EnsLLM’s accuracy but introduce notable computational cost. In future work, we plan to explore strategies for more efficient ensemble execution, including selective model invocation, partial ensembling, or model pruning. We also aim to investigate system-level solutions for running multiple LLMs concurrently with minimal resource overhead, to make the approach more practical for real-world deployment.
Table 4: Performance of EnsLLM compared to standalone LLMs on HumanEval-Java(pass@1). The accuracy provided in the braces is achievable accuracy.
LLM Accuracy (%)
GPT-4o 90
GPT-4 84
GPT-3.5 78
CoPilot 86
OpenChat (F) 70
Llama 3.2 (F) 72
Gemini 86
Qwen2 (F) 72
Codestral (F) 68
Gemma 2 (F) 62
DeepSeekCoder (F) 60
CodeLlama (F) 64
DolphinCoder (F) 48
EnsLLM (All) 90 (90)
5.3 Case Study on HumanEval-Java
To evaluate if EnsLLM works for other languages, we conducted a case study on the HumanEval-Java [1] benchmark, which closely follows the structure and semantics of HumanEval but is written in Java. For syntactic and semantic similarity, we reused CodeBLEU, which supports multi-language analysis without modification. However, since CrossHair is Python-specific, we choose an alternative approach to approximate behavioral similarity for Java programs.
We employed Kex [2], a symbolic execution-based test generation framework for Java, as a substitute for CrossHair in the Java setting. Kex performs path-sensitive analysis by symbolically executing Java bytecode to explore multiple execution paths and automatically synthesize input values that trigger them. For each candidate program $P_{i}$, we used Kex to generate a suite of test cases that exercise its behavior over a wide range of symbolic inputs. These tests aim to reflect the program’s logic and edge cases without requiring manually written specifications.
To compute behavioral similarity between two programs $P_{i}$ and $P_{j}$, we leveraged these generated test cases in a cross-execution strategy. Specifically, for every program pair $(P_{i},P_{j})$, we executed the test suite synthesized for $P_{i}$ on the implementation of $P_{j}$. We then counted the number of test cases that failed when run against $P_{j}$—i.e., those that produced exceptions, assertion failures, or incorrect return values. This count was denoted as $cex(P_{i},P_{j})$ and treated as a proxy for behavioral divergence from $P_{i}$’s semantics.
To convert this into a normalized similarity measure, we applied the same formulation used in Equation (1), capping the counterexamples at $n=6$ to avoid extreme penalty due to overly sensitive or noisy test cases.
Due to the significant effort required to integrate this new workflow, we report here on a preliminary evaluation on the first 50 problems from the HumanEval-Java dataset. We exclude CodeBERT from this experiment as its performance is hindered by the prompt structure used in HumanEval-Java. Unlike benchmarks like HumanEval or LiveCodeBench, HumanEval-Java lacks well-researched prompt engineering, leading CodeBERT, which is a masked language model, to return the input without generating meaningful method bodies.
As shown in Table 4, EnsLLM achieved an accuracy of 90%, which matches the achievable accuracy. Although the accuracy of the best-performing standalone model in our experiment, GPT-4o, is also 90%, EnsLLM selected the correct solution in every case where a correct solution was available. These findings suggest that the core principles of our ensemble method generalize well across programming languages, provided that compatible analysis tools are available or can be adapted.
@@ 5.4 Threats to Validity
For external validity, our evaluation is limited to HumanEval and LiveCodeBench, which may not fully represent complex, real-world programming challenges. However, these two benchmarks are widely recognized in code generation research and have been used in several similar studies to evaluate the functional correctness of LLM-generated programs. Their extensive use ensures a standardized comparison with existing approaches, mitigating the threat to generalizability. A potential threat to external validity lies in the selection of the 14 LLMs used in our ensemble. While we included a mix of open-source and proprietary models with varied sizes and tuning, the selection was not based on a formal strategy. This may limit the generalizability of our results, particularly regarding robustness to correlated errors. Future work will explore principled selection methods based on model diversity and behavioral variance to strengthen ensemble design. Another external validity concern is the potential presence of the solution of the datasets in the pretraining sets of some LLMs. While we cannot verify the training data of all models, any such overlap could inflate performance. Although our focus is on the ensemble mechanism rather than generation quality, this remains a limitation.
@@ 6 Related Work
LLMs have also been explored in broader software engineering tasks, including automated bug detection, test case generation, code summarization, and security vulnerability analysis [27, 23]. For instance, they can assist in writing unit tests by generating test cases from function signatures, improving test coverage while reducing developer effort [mueller2023automated]. Additionally, LLMs are being integrated into code refactoring tools, where they suggest optimizations and enhance code readability [20]. Beyond these applications, LLMs are transforming software engineering workflows, influencing how software is developed [25, 34], enhancing developer productivity [16, 47], and assisting in security analysis [39]. However, challenges persist in ensuring correctness, maintainability, and security, requiring further refinement and validation before LLM-generated code can be reliably integrated into real-world software systems.
Test-based selection has been explored to improve code generation. CodeT [13] uses LLM-generated tests to rank candidate programs from the same model based on execution results and agreement. Fakhoury et al. [19] similarly apply unit test feedback in an interactive loop to guide code refinement. While conceptually aligned with our goal of functional selection, these methods rely on a single model, limiting candidate diversity. On the other hand, EnsLLM leverages multiple LLMs, reducing correlated errors and enabling more robust selection.
The mixture of experts (MoE) framework [28] combines multiple specialized models with a gating network to select the best expert for each input, using Expectation-Maximization for training. Fedus et al. [21] scaled MoE to trillion-parameter Switch Transformers with sparse activation for efficient language processing. Despite these advancements, computational efficiency remains a challenge when using multiple models for selection and refinement. Also, EnsLLM is the first approach that ensembles multiple LLMs for code generation.